# What Layers When: Learning to Skip Compute in LLMs with Residual Gates

### v1

@@ 1 Introduction
    We propose GateSkip, a residual gating mechanism that enables token-wise layer skipping with smooth training and deterministic hard decisions at inference. [^1] The code is available at https://github.com/Thiggel/GateSkip
@@ 2 Related Work
A growing body of work has addressed the inefficiency of ever‑larger decoder‑only Transformers by dynamically adapting computation on a per‑token or per‑sequence basis. Early efforts in layer skipping repurpose sparse routing from Mixture‑of‑Experts (MoE) to decide whether to execute each layer at all, yielding substantial compute savings without restarting training from scratch. Mixture‑of‑Depths (MoD) injects a router at every transformer layer to skip unimportant layers for each token (Raposo et al., 2024), while follow‑up methods introduce soft token budgets (Zeng et al., 2023), sequence‑level skipping (Wang et al., 2023), and frozen‑backbone router fine‑tuning (He et al., 2024). However, discrete routers can be unstable (Zoph et al., 2022; Fedus et al., 2022; Puigcerver et al., 2024; Panda et al., 2025), and most require training from scratch.
In contrast to hard top-k routers with load-balancing losses, GateSkip uses fully differentiable residual gating, avoiding discrete routing during training while still yielding hard skips at test time.
In contrast, early exiting methods terminate inference once intermediate representations are deemed confident enough. Pioneered in encoder‑only BERT models via entropy or agreement thresholds (Xin et al., 2020; Zhang et al., 2022; Zhou et al., 2020; Liu et al., 2020), this concept has been extended to encoder–decoder and decoder‑only settings by supervising every layer with an auxiliary language modeling objective and exiting based on heuristic confidence measures (Tang et al., 2023; Schuster et al., 2022; Elbayad et al., 2020; Liu et al., 2021; Bae et al., 2023; Elhoushi et al., 2024; Del Corro et al., 2023). Yet these approaches fundamentally alter pretrained hidden representations through connector modules or self‑distillation (Bajpai & Hanawal, 2024a; b; Ji et al., 2023).
Where early exit supervises intermediate LM heads and exits on confidence thresholds, GateSkip avoids auxiliary heads and maintains pretrained hidden spaces by training gates to compress post-module outputs.
@@ 3.4 Implementation Details
The added gates introduce negligible parameter overhead to the model, e.g. 0.004% for separate gates with scalar output, and 4% for separate gates with hidden-state sized output on Llama-3.2-1b. Table 13 in Appendix G shows parameter overhead for each of the variants of GateSkip.
@@ 4 Experiments
Llama-1b 31.3 - - - - - 45.0 - - - -
Llama-1b (random skipping) - 9.7 1.9 3.0 1.5 1.1 - 32.0 31.5 31.1 31.0
CALM (hidden state saturation) 0.6 0.7 0.6 0.4 0.5 0.6 34.9 32.7 31.9 30.7 31.1
CALM (softmax) 0.8 0.9 1.0 0.3 0.2 0.4 34.9 32.3 30.8 31.1 31.1
FREE (hidden state saturation) 11.9 11.9 11.9 11.9 11.9 11.9 38.6 38.6 38.6 38.6 38.6
FREE (softmax) 11.9 11.9 11.9 11.9 11.9 11.9 38.6 38.6 38.6 38.6 38.6
LayerSkip 13.1 13.1 13.1 13.1 13.1 13.1 38.3 38.3 38.3 38.3 38.3
MoD (router-tuned) 20.1 17.3 14.5 7.7 5.8 4.0 40.6 35.4 34.5 31.3 29.6
SkipLayer 2.2 1.4 0.7 0.0 0.0 0.0 31.2 30.7 30.2 30.2 30.3
GateSkip 26.8 25.5 24.2 23.2 23.6 19.8 44.3 37.8 32.7 30.8 31.2
@@ 4.1 Experimental Setup
We primarily evaluate our method on Llama-3.2-1b, while also experimenting with Llama-3.2-3b, Llama-3.1-8b, and Gemma-2-2b to assess scalability and architecture independence. For all experiments, we fine-tune the pretrained backbone to train the gates while simultaneously adapting the model to task templates for easier downstream answer extraction. We set the sparsity loss weight $\lambda=0.1$ and decay the token budget from 100% to 80% during training. Training employs the AdamW optimizer (Loshchilov & Hutter, 2017). A full list of hyperparameters can be found in Appendix L. All libraries and their respective versions used for our experiments are listed in Appendix M. Instructions for code access can be found in Appendix I.
CSQA
GSM8K (Gen.)
MMLU
Stem
HellaSwag CSQA PIQA
Open-
BookQA
Wino-
Grande
Llama-1b (random skipping) 6.0 0.0 24.9 29.0 19.0 49.3 17.3 52.5
CALM (hidden state saturation) 0.2 0.7 23.4 29.4 18.9 56.2 15.2 32.7
CALM (softmax) 0.0 0.8 25.7 29.3 19.3 53.2 15.0 51.2
MoD (router-tuned) 8.8 6.5 26.5 35.0 18.5 58.8 20.3 53.5
GateSkip 36.7 9.7 26.0 38.5 21.2 66.5 21.3 53.0
@@
We fine-tune on the train sets of CommonsenseQA (Talmor et al., 2019) and GSM8K (Cobbe et al., 2021) questions with chain-of-thought traces generated by Nemotron-70B (Anonymous, 2024; Reasoning, 2024; Wang et al., 2024), masking the loss on the question portion so that the model learns both reasoning and answer extraction. For evaluation we measure zero-shot accuracy on the GSM8K and CommonsenseQA test sets using the same prompt template. We sweep the inference budget $\hat{b}$ over $\{1.00,0.95,0.90,0.85,0.80,0.75\}$, corresponding to compute-savings of $\{0\%,5\%,10\%,15\%,20\%,25\%\}$; since the realized savings sometimes fall between these targets, we linearly interpolate the measured accuracies to report performance at the exact percentages listed above. Non-interpolated results are reported in Appendix K.
@@
Each full run requires $\approx 5$ GPU-hours on an H100-80GB, so we perform one seed per configuration and estimate uncertainty by bootstrapping (100 000 resamples in LM-Eval-Harness), yielding a standard deviation $\leq 3\%$ on every benchmark. Standard deviations and raw results are reported in Appendix K. All baselines (random skipping, MoD router-tuning, CALM variants) were trained and evaluated with the identical data splits, hyperparameters, inference budgets, and interpolation procedure described above, ensuring a fair comparison.
@@ 4.2 Comparison to Baseline
Table 3: GateSkip on Llama Instruct.
@@
Table 4: GateSkip on out-of-domain generative tasks.
Table 1 presents averaged accuracies across multiple compute‐savings levels, while Table 2 reports performance at exactly 15% saved compute. Random skipping collapses generative accuracy to under 10% even at modest budgets (5–10% savings), while GateSkip retains the majority of performance. Full results can be seen in Appendix K & J.
@@ 4.4 Component Ablations
Table 5: Ablation of GateSkip’s design choices on Llama-3.2-1b.
@@
Table 6: GateSkip on different model sizes and architectures.
@@
To test scalability, we applied GateSkip to larger Llama models (3b and 8b) and observed consistent performance patterns (see Table 6). Full results and plots are shown in Appendix K and J. The results reveal that for larger architectures, the model is capable of skipping increasingly more tokens without decreasing performance. For instance, Llama-3.2-3B with GateSkip can save 37.3% computation while retaining 91.5% of its baseline GSM8K (Gen.) performance and 87.3% of its baseline CSQA (Gen.) performance. Moreover, comparisons between Llama and Gemma architectures reveal that the compute-accuracy trade-off generalizes across both model families. The instruction-tuned and LAMBADA results mirror these trends: larger or instruction-adapted backbones benefit more from budgeted token selection, maintaining strong generation quality where random or uniform skipping fails, and confirming that GateSkip’s gains persist across usage styles (chat/instruction), lengths, and domains.

### v3

@@ 1 Introduction
    We propose GateSkip, a residual gating mechanism that enables token-wise layer skipping with smooth training and deterministic hard decisions at inference.
@@ 2 Related Work
A growing body of work has addressed the inefficiency of ever‑larger decoder‑only Transformers by dynamically adapting computation on a per‑token or per‑sequence basis. Early efforts in layer skipping repurpose sparse routing from Mixture‑of‑Experts (MoE) to decide whether to execute each layer at all, yielding substantial compute savings without restarting training from scratch. Mixture‑of‑Depths (MoD) injects a router at every transformer layer to skip unimportant layers for each token (Raposo et al., 2024), while follow‑up methods introduce soft token budgets (Zeng et al., 2023), sequence‑level skipping (Wang et al., 2023), and frozen‑backbone router fine‑tuning (He et al., 2024). However, discrete routers can be unstable (Zoph et al., 2022; Fedus et al., 2022; Puigcerver et al., 2024; Panda et al., 2025), and most require training from scratch. In contrast to hard top-k routers with load-balancing losses, GateSkip uses fully differentiable residual gating, avoiding discrete routing during training while still yielding hard skips at test time.
In contrast, early exiting methods terminate inference once intermediate representations are deemed confident enough. Pioneered in encoder‑only BERT models via entropy or agreement thresholds (Xin et al., 2020; Zhang et al., 2022; Zhou et al., 2020; Liu et al., 2020), this concept has been extended to encoder–decoder and decoder‑only settings by supervising every layer with an auxiliary language modeling objective and exiting based on heuristic confidence measures (Tang et al., 2023; Schuster et al., 2022; Elbayad et al., 2020; Liu et al., 2021; Bae et al., 2023; Elhoushi et al., 2024; Del Corro et al., 2023). Yet these approaches fundamentally alter pretrained hidden representations through connector modules or self‑distillation (Bajpai & Hanawal, 2024a; b; Ji et al., 2023). Where early exit supervises intermediate LM heads and exits on confidence thresholds, GateSkip avoids auxiliary heads and maintains pretrained hidden spaces by training gates to compress post-module outputs.
@@ 3.4 Implementation Details
The added gates introduce negligible parameter overhead to the model, e.g. 0.004% for separate gates with scalar output, and 4% for separate gates with hidden-state sized output on Llama-3.2-1b. Table 13 in Appendix H shows parameter overhead for each of the variants of GateSkip.
@@ 4 Experiments
Llama-1b $\mathbf{30.97_{\scriptscriptstyle\pm 1.44}}$ - - - - - $49.12_{\scriptscriptstyle\pm 0.05}$ - - - -
Llama-1b (random skipping) - $10.30_{\scriptscriptstyle\pm 0.54}$ $2.20_{\scriptscriptstyle\pm 0.36}$ $1.67_{\scriptscriptstyle\pm 0.25}$ $1.23_{\scriptscriptstyle\pm 0.26}$ $0.67_{\scriptscriptstyle\pm 0.12}$ - $25.58_{\scriptscriptstyle\pm 0.25}$ $23.62_{\scriptscriptstyle\pm 0.27}$ $23.36_{\scriptscriptstyle\pm 0.20}$ $23.65_{\scriptscriptstyle\pm 0.05}$
CALM (hidden state saturation) $3.43_{\scriptscriptstyle\pm 0.49}$ $3.43_{\scriptscriptstyle\pm 0.49}$ $3.43_{\scriptscriptstyle\pm 0.49}$ $3.43_{\scriptscriptstyle\pm 0.49}$ $3.43_{\scriptscriptstyle\pm 0.49}$ $3.43_{\scriptscriptstyle\pm 0.49}$ $30.73_{\scriptscriptstyle\pm 0.00}$ $30.73_{\scriptscriptstyle\pm 0.00}$ $30.73_{\scriptscriptstyle\pm 0.00}$ $30.73_{\scriptscriptstyle\pm 0.00}$ $30.73_{\scriptscriptstyle\pm 0.00}$
CALM (softmax) $3.15_{\scriptscriptstyle\pm 0.35}$ $3.15_{\scriptscriptstyle\pm 0.35}$ $3.15_{\scriptscriptstyle\pm 0.35}$ $3.15_{\scriptscriptstyle\pm 0.35}$ $3.15_{\scriptscriptstyle\pm 0.35}$ $3.15_{\scriptscriptstyle\pm 0.35}$ $30.73_{\scriptscriptstyle\pm 0.00}$ $30.73_{\scriptscriptstyle\pm 0.00}$ $30.73_{\scriptscriptstyle\pm 0.00}$ $30.73_{\scriptscriptstyle\pm 0.00}$ $30.73_{\scriptscriptstyle\pm 0.00}$
FREE (hidden state saturation) $1.57_{\scriptscriptstyle\pm 0.63} 11$$57_{\scriptscriptstyle\pm 0.63} 11.$$57_{\scriptscriptstyle\pm 0.63} 11.$$57_{\scriptscriptstyle\pm 0.63} 11.$$57_{\scriptscriptstyle\pm 0.63} 11.$$1.57_{\scriptscriptstyle\pm 0.63}$ $36.02_{\scriptscriptstyle\pm 0.01}$ $36.02_{\scriptscriptstyle\pm 0.01}$ $36.02_{\scriptscriptstyle\pm 0.01}$ $36.02_{\scriptscriptstyle\pm 0.01}$ $36.02_{\scriptscriptstyle\pm 0.01}$
FREE (softmax) $10.70_{\scriptscriptstyle\pm 0.00}$ $10.70_{\scriptscriptstyle\pm 0.00}$ $10.70_{\scriptscriptstyle\pm 0.00}$ $10.70_{\scriptscriptstyle\pm 0.00}$ $10.70_{\scriptscriptstyle\pm 0.00}$ $10.70_{\scriptscriptstyle\pm 0.00}$ $36.02_{\scriptscriptstyle\pm 0.01}$ $36.02_{\scriptscriptstyle\pm 0.01}$ $36.02_{\scriptscriptstyle\pm 0.01}$ $36.02_{\scriptscriptstyle\pm 0.01}$ $36.02_{\scriptscriptstyle\pm 0.01}$
LayerSkip $10.65_{\scriptscriptstyle\pm 1.55}$ $10.65_{\scriptscriptstyle\pm 1.55}$ $10.65_{\scriptscriptstyle\pm 1.55}$ $10.65_{\scriptscriptstyle\pm 1.55}$ $10.65_{\scriptscriptstyle\pm 1.55}$ $10.65_{\scriptscriptstyle\pm 1.55}$ $38.25_{\scriptscriptstyle\pm 0.02}$ $38.25_{\scriptscriptstyle\pm 0.02}$ $\mathbf{38.25_{\scriptscriptstyle\pm 0.02}}$ $\mathbf{38.25_{\scriptscriptstyle\pm 0.02}}$ $\mathbf{38.25_{\scriptscriptstyle\pm 0.02}}$
MoD (router-tuned) $_{\scriptscriptstyle\pm 0.00} 0.00±0.00$$_{\scriptscriptstyle\pm 0.00}$ $_{\scriptscriptstyle\pm 0.00} 0.00±0.00$$_{\scriptscriptstyle\pm 0.00}$ $_{\scriptscriptstyle\pm 0.00} 0.00±0.00$$_{\scriptscriptstyle\pm 0.00}$ $\mathbf{51.34_{\scriptscriptstyle\pm 0.05}}$ $3.68_{\scriptscriptstyle\pm 0.03} 22$$98_{\scriptscriptstyle\pm 0.23} 22.$$94_{\scriptscriptstyle\pm 0.09} 22.$$2.86_{\scriptscriptstyle\pm 0.04}$
MoD $20.83_{\scriptscriptstyle\pm 5.22}$ $9.45_{\scriptscriptstyle\pm 4.63}$ $4.90_{\scriptscriptstyle\pm 2.92}$ $96_{\scriptscriptstyle\pm 1.67} 3.3$$37_{\scriptscriptstyle\pm 1.33}$ $2.91_{\scriptscriptstyle\pm 1.33}$ $44.18_{\scriptscriptstyle\pm 0.66}$ $31.88_{\scriptscriptstyle\pm 1.48}$ $29.33_{\scriptscriptstyle\pm 2.42}$ $26.72_{\scriptscriptstyle\pm 0.39}$ $26.00_{\scriptscriptstyle\pm 0.58}$
SkipLayer $0.70_{\scriptscriptstyle\pm 0.60}$ $24_{\scriptscriptstyle\pm 0.17} 0.0$$4_{\scriptscriptstyle\pm 0.00} 0.0$$5_{\scriptscriptstyle\pm 0.05} 0.0$$5_{\scriptscriptstyle\pm 0.05} 0.0$$04_{\scriptscriptstyle\pm 0.04}$ $32.00_{\scriptscriptstyle\pm 1.46}$ $23.03_{\scriptscriptstyle\pm 0.12}$ $23.60_{\scriptscriptstyle\pm 0.44}$ $23.67_{\scriptscriptstyle\pm 0.47}$ $23.41_{\scriptscriptstyle\pm 0.09}$
GateSkip (ours) $23.53_{\scriptscriptstyle\pm 1.28}$ $\mathbf{23.05_{\scriptscriptstyle\pm 1.24}}$ $\mathbf{22.57_{\scriptscriptstyle\pm 1.24}}$ $\mathbf{22.14_{\scriptscriptstyle\pm 1.36}}$ $\mathbf{20.37_{\scriptscriptstyle\pm 1.18}}$ $\mathbf{17.67_{\scriptscriptstyle\pm 0.74}}$ $47.35_{\scriptscriptstyle\pm 0.22}$ $\mathbf{38.86_{\scriptscriptstyle\pm 0.45}}$ $31.74_{\scriptscriptstyle\pm 2.04}$ $27.93_{\scriptscriptstyle\pm 0.47}$ $26.49_{\scriptscriptstyle\pm 0.57}$
@@ 4.1 Experimental Setup
We primarily evaluate our method on Llama-3.2-1b, while also experimenting with Llama-3.2-3b, Llama-3.1-8b, and Gemma-2-2b to assess scalability and architecture independence. For all experiments, we fine-tune the pretrained backbone to train the gates while simultaneously adapting the model to task templates for easier downstream answer extraction. We set the sparsity loss weight $\lambda=0.1$ and decay the token budget from 100% to 80% during training. Training employs the AdamW optimizer (Loshchilov & Hutter, 2017). A full list of hyperparameters can be found in Appendix K. All libraries and their respective versions used for our experiments are listed in Appendix L. Instructions for code access can be found in Appendix J.
@@
CSQA (Gen.) GSM8K (Gen.) MMLU Stem HellaSwag CSQA PIQA Open- BookQA Wino- Grande
Llama-1b (random skipping) $3.27_{\scriptscriptstyle\pm 0.53}$ $0.07_{\scriptscriptstyle\pm 0.09}$ $24.00_{\scriptscriptstyle\pm 0.26}$ $30.87_{\scriptscriptstyle\pm 0.57}$ $20.27_{\scriptscriptstyle\pm 1.09}$ $53.67_{\scriptscriptstyle\pm 0.94}$ $15.74_{\scriptscriptstyle\pm 0.66}$ $51.13_{\scriptscriptstyle\pm 0.34}$
CALM (hidden state saturation) $5.40_{\scriptscriptstyle\pm 1.23}$ $1.47_{\scriptscriptstyle\pm 0.25}$ $21.85_{\scriptscriptstyle\pm 0.00}$ $34.00_{\scriptscriptstyle\pm 0.00}$ $19.00_{\scriptscriptstyle\pm 0.00}$ $61.60_{\scriptscriptstyle\pm 0.00}$ $16.20_{\scriptscriptstyle\pm 0.00}$ $50.60_{\scriptscriptstyle\pm 0.00}$
CALM (softmax) $4.70_{\scriptscriptstyle\pm 0.90}$ $1.60_{\scriptscriptstyle\pm 0.20}$ $21.85_{\scriptscriptstyle\pm 0.00}$ $34.00_{\scriptscriptstyle\pm 0.00}$ $19.00_{\scriptscriptstyle\pm 0.00}$ $61.60_{\scriptscriptstyle\pm 0.00}$ $16.20_{\scriptscriptstyle\pm 0.00}$ $50.60_{\scriptscriptstyle\pm 0.00}$
FREE (hidden state saturation) $17.07_{\scriptscriptstyle\pm 0.57}$ $6.07_{\scriptscriptstyle\pm 0.75}$ $21.26_{\scriptscriptstyle\pm 0.01}$ $41.33_{\scriptscriptstyle\pm 0.09}$ $19.40_{\scriptscriptstyle\pm 0.00}$ $70.93_{\scriptscriptstyle\pm 0.09}$ $21.87_{\scriptscriptstyle\pm 0.09}$ $50.47_{\scriptscriptstyle\pm 0.09}$
FREE (softmax) $16.40_{\scriptscriptstyle\pm 0.00}$ $5.00_{\scriptscriptstyle\pm 0.00}$ $21.27_{\scriptscriptstyle\pm 0.02}$ $41.30_{\scriptscriptstyle\pm 0.10}$ $19.40_{\scriptscriptstyle\pm 0.00}$ $71.00_{\scriptscriptstyle\pm 0.00}$ $21.90_{\scriptscriptstyle\pm 0.10}$ $50.30_{\scriptscriptstyle\pm 0.10}$
LayerSkip $16.40_{\scriptscriptstyle\pm 1.80}$ $4.90_{\scriptscriptstyle\pm 1.30}$ $21.31_{\scriptscriptstyle\pm 0.00}$ $\mathbf{42.40_{\scriptscriptstyle\pm 0.00}}$ $19.00_{\scriptscriptstyle\pm 0.00}$ $\mathbf{72.33_{\scriptscriptstyle\pm 0.09}}$ $20.80_{\scriptscriptstyle\pm 0.00}$ $\mathbf{56.33_{\scriptscriptstyle\pm 0.09}}$
MoD (router-tuned) $_{\scriptscriptstyle\pm 0.00} 0.00±0.00$$_{\scriptscriptstyle\pm 0.00}$ $23.09_{\scriptscriptstyle\pm 0.21}$ $28.87_{\scriptscriptstyle\pm 0.35}$ $19.92_{\scriptscriptstyle\pm 0.24}$ $53.60_{\scriptscriptstyle\pm 0.41}$ $13.81_{\scriptscriptstyle\pm 0.54}$ $49.08_{\scriptscriptstyle\pm 1.36}$
MoD $6.14_{\scriptscriptstyle\pm 3.06}$ $1.77_{\scriptscriptstyle\pm 0.29}$ $26.00_{\scriptscriptstyle\pm 0.66}$ $38.21_{\scriptscriptstyle\pm 1.55}$ $20.86_{\scriptscriptstyle\pm 0.94}$ $59.88_{\scriptscriptstyle\pm 2.21}$ $18.58_{\scriptscriptstyle\pm 1.36}$ $52.15_{\scriptscriptstyle\pm 0.73}$
SkipLayer $10_{\scriptscriptstyle\pm 0.10} 0.0$$00_{\scriptscriptstyle\pm 0.00}$ $21.82_{\scriptscriptstyle\pm 0.42}$ $27.76_{\scriptscriptstyle\pm 0.80}$ $19.06_{\scriptscriptstyle\pm 1.67}$ $50.51_{\scriptscriptstyle\pm 0.50}$ $16.40_{\scriptscriptstyle\pm 1.23}$ $48.72_{\scriptscriptstyle\pm 1.11}$
GateSkip (ours) $\mathbf{35.25_{\scriptscriptstyle\pm 1.81}}$ $\mathbf{9.03_{\scriptscriptstyle\pm 1.31}}$ $\mathbf{30.86_{\scriptscriptstyle\pm 1.18}}$ $39.05_{\scriptscriptstyle\pm 1.25}$ $\mathbf{36.16_{\scriptscriptstyle\pm 1.23}}$ $70.45_{\scriptscriptstyle\pm 0.60}$ $\mathbf{22.69_{\scriptscriptstyle\pm 0.43}}$ $52.66_{\scriptscriptstyle\pm 0.82}$
@@
We fine-tune on the train sets of CommonsenseQA (Talmor et al., 2019) and GSM8K (Cobbe et al., 2021) questions with chain-of-thought traces generated by Nemotron-70B (Anonymous, 2024; Reasoning, 2024; Wang et al., 2024), masking the loss on the question portion so that the model learns both reasoning and answer extraction. For evaluation we measure zero-shot accuracy on the GSM8K and CommonsenseQA test sets using the same prompt template. We sweep the inference budget $\hat{b}$ over $\{1.00,0.95,0.90,0.85,0.80,0.75\}$, corresponding to compute-savings of $\{0\%,5\%,10\%,15\%,20\%,25\%\}$; since the realized savings sometimes fall between these targets, we linearly interpolate the measured accuracies to report performance at the exact percentages listed above.
@@
We perform five seeds per configuration, measuring mean and standard deviation on each metric. All baselines (random skipping, MoD router-tuning, CALM variants) are trained and evaluated with the identical data splits, hyperparameters, inference budgets, and interpolation procedure described above, ensuring a fair comparison.
@@ 4.2 Comparison to Baseline
Table 4: GateSkip on out-of-domain generative tasks.
@@
Table 1 presents averaged accuracies across multiple compute‐savings levels, while Table 2 reports performance at exactly 15% saved compute. Random skipping collapses generative accuracy to under 10% even at modest budgets (5–10% savings), while GateSkip retains the majority of performance.
@@ 4.4 Component Ablations
Table 5: Ablation of GateSkip’s design choices on Llama-3.2-1b.
@@
Gate Loss
L1 23.0 22.6 22.2 21.4 18.7 17.0 47.5 37.6 29.2 27.1 26.3
KL-div. 36.0 18.1 11.9 9.8 7.9 5.3 45.6 26.6 24.5 23.0 23.0
L2 (default) 26.8 25.5 24.2 23.2 23.6 19.8 44.3 37.8 32.7 30.8 31.2
Frozen Backbone
Frozen B. 14.9 14.1 13.3 12.7 12.5 12.6 45.6 37.5 26.7 24.0 23.9
Unfrozen B. (default) 26.8 25.5 24.2 23.2 23.6 19.8 44.3 37.8 32.7 30.8 31.2
@@
Gate Loss Type. Replacing our L2-loss on gate activations with an L1 loss mostly underperforms, whereas it achieves a higher loglikelihood accuracy at 0% compute savings. Furthermore, we tested a layer-wise KL divergence regularizer that matches the average gate activation per layer to a target budget, encouraging control of the overall fraction of retained tokens rather than only shrinking individual gates:
$\mathcal{L}_{\text{layer-KL}}=\sum_{\ell}\mathrm{KL}\!\left(\mathrm{Bern}(\bar{g}_{\ell})\;\|\;\mathrm{Bern}(b_{\text{train}})\right),\text{where}\ \bar{g}_{\ell}=\frac{1}{BSH}\sum_{b,i,k}g_{\ell}(h_{\ell})_{b,i,k}.$
While this loss clearly outperforms our L2 and L1 variants at no compute savings in the generative settings, it performs poorly at even low compute savings.
Frozen Backbone. To test the extent to which the backbone adapts, we froze a backbone finetuned on our dataset (for fair comparison) and trained gates on the frozen backbone. In this frozen setting, downstream accuracy substantially decreases. This gives evidence that backbone adaptation plays a major role in the skipping task: the backbone may by itself not reveal all information needed for skipping, necessitation adaptation so that it saves relevance cues in the hidden states which in turn condition the gates.
Table 6: GateSkip on different model sizes and architectures.
@@
To test scalability, we applied GateSkip to larger Llama models (3b and 8b) and observed consistent performance patterns (see Table 6). The results reveal that for larger architectures, the model is capable of skipping increasingly more tokens without decreasing performance. For instance, Llama-3.2-3B with GateSkip can save 37.3% computation while retaining 91.5% of its baseline GSM8K (Gen.) performance and 87.3% of its baseline CSQA (Gen.) performance. Moreover, comparisons between Llama and Gemma architectures reveal that the compute-accuracy trade-off generalizes across both model families. The instruction-tuned and LAMBADA results mirror these trends: larger or instruction-adapted backbones benefit more from budgeted token selection, maintaining strong generation quality where random or uniform skipping fails, and confirming that GateSkip’s gains persist across usage styles (chat/instruction), lengths, and domains.